# Navigate the Unknown: Enhancing LLM Reasoning with Intrinsic Motivation Guided Exploration

### v5

@@ 2.2.2 Error-conditioned Reward Allocation
Despite the benefits of trajectory-aware exploration rewards, the large action space in LLM reasoning introduces training challenges. Naive exploration strategies become infeasible due to unstable and sparse rewards and suboptimal exploration efficiency. Moreover, uniform exploration rewards across correct and incorrect samples reduce exploration efficiency and induce instability in correct trajectories. To enable more effective exploration, we prioritize samples with residual errors to facilitate diverse discoveries while efficiently improving overall performance. Thus, i-MENTOR employs an error-conditioned reward allocation mechanism across sampling sequences. Specifically, a selection function $I_{a\neq o}$ is designed to allocate exploration rewards exclusively to incorrect samples:
@@ 2.2.3 Advantage-preserving Integration
Since the intensity of the exploration behavior varies at different stages of training, directly combining exploration rewards with outcome-based rewards creates conflicting learning signals across training stages. For PPO, this injects noise into value function estimation. For GRPO, exploration rewards may invert originally positive advantage signs after group normalization, which is a critical issue since exploration rewards should never penalize samples. To resolve this conflict and ensure seamless integration with established RL algorithms like PPO and GRPO, i-MENTOR applies exploration rewards after advantage computation and possible value estimation through:
@@
where $\hat{A}_{\text{old}}$ denotes the original advantage from outcome-based rewards. For PPO, we apply the exploration reward to the last advantage token. For GRPO, we apply it to each rollout advantage. This design preserves two key properties: (1) The outcome-based advantage maintains its original mean and variance statistics, ensuring stable policy updates; (2) Exploration rewards $R^{\star}(q,o)$ provide trajectory-level guidance without distorting value estimation. The decoupled formulation prevents exploration rewards from conflicting with outcome-based advantage normalization in GRPO or perturbing value function update in PPO, enabling harmonious integration of both reward components.
By applying trajectory-aware exploration rewards with error-conditioned reward allocation and advantage-preserving integration, i-MENTOR efficiently enhances LLM reasoning by guiding the model toward diverse possible responses during training, while significantly improving overall reasoning ability. The complete algorithm is detailed in Appendix B.
@@ 3 Experiments
Here we evaluate i-MENTOR on three datasets to investigate the following research questions:
  • Q1: How does i-MENTOR enhance LLM reasoning performance?
  • Q2: How does i-MENTOR generalize across models and data distributions?
In the following subsections, we begin by outlining the experimental setup, followed by a systematic results analysis addressing our core questions. Training statistics are visualized in Figure 3. Appendix A provides an expanded case study for Figure 2, while implementation specifics–including pseudo code (Appendix B), data statistics (Appendix C), and configuration details (Appendix D)–are deferred for completeness. Beyond core experiments, Appendix E validates i-MENTOR’s superiority over basic exploration techniques. We further include an ablation study (Appendix F) quantifying submodule contributions and a sensitivity analysis (Appendix G) examining hyperparameter impacts on performance.
@@ 3.1.1 Dataset
We conduct experiments on three public datasets, i.e., GSM8K(Cobbe et al., 2021), Countdown-34 and its harder version Countdown-4(Gandhi et al., 2024) to validate i-MENTOR’s effectiveness against vanilla PPO and GRPO algorithms. For computational efficiency, we use a subset of the complete dataset of Countdown-34 and Countdown-4 for training. The detailed dataset statistic is shown in Appendix C. Additionally, we also provide detailed experiments on the AIME 2024 dataset between GRPO and i-MENTOR-GRPO to further validate i-MENTOR’s effectiveness with multiple outputs in Table 2.
3.1.2 Evaluation Protocol
Network structure of i-MENTOR: The predictor and target networks in i-MENTOR share the same architecture, which is dataset-agnostic and LLM-independent. For computational efficiency, we simply apply a lightweight predictor and target network structure for i-MENTOR (embedding size=16=16, three FFN layers with neurons $[16,8,1]$), ensuring negligible additional training time. To make a fairer comparison, we apply a set of fixed training hyperparameters to all datasets.
3.2 Main Result (Q1)
The main results are presented in Table 1. Figure 4 highlights key training details via some of the training trajectories, showcasing the performance improvements achieved during training.
Table 1: Average accuracy. i-MENTOR-PPO/GRPO denotes the implementation of i-MENTOR on PPO/GRPO. “Improve” indicates relative improvement.
Dataset PPO i-MENTOR-PPO Improve GRPO i-MENTOR-GRPO Improve
GSM8K 0.8051 0.8169 1.47% 0.8082 0.8251 2.09%
Countdown-34 0.5526 0.5924 7.2% 0.6711 0.7132 6.27%
Countdown-4 0.3307 0.3812 15.27% 0.3872 0.4739 22.39%
@@ 3.2 Main Result (Q1)
From Table 1,  2 and Figure 4 we could conclude that:
  • GRPO outperforms PPO on all datasets, illustrating superior training effects brought by group-wise sampling and reward normalization. Both i-MENTOR-PPO and i-MENTOR-GRPO outperform standard RL approaches, demonstrating that i-MENTOR effectively guides and enhances the updates of LLM reasoning via exploration behavior during the RL training process. By encouraging the exploration of new responses rather than merely fitting to outcome-based reward through sampling behaviors, i-MENTOR enables LLMs to explore more diverse potential inference paths during training and avoids LLMs getting trapped in local optimal solutions.
  • A key challenge in improving LLM reasoning lies in handling difficult samples, which typically yield near-zero rewards that hinder parameter updates. Our experiments reveal an interesting pattern: i-MENTOR’s improvements are most significant on Countdown-4, followed by Countdown-34 and GSM8K. This progression aligns with the relative difficulty levels of these datasets, where Countdown-4 > Countdown-34 > GSM8K. The results suggest that i-MENTOR’s dense exploration rewards specifically help models overcome learning barriers in challenging samples. This phenomenon represents a further contribution of our study: it originates from RLVR’s core distinction from traditional RL–leveraging a pre-trained LLM instead of a randomly initialized model. Consequently, exploration yields significantly greater benefits when addressing complex problems.
  • i-MENTOR achieves greater performance gains with GRPO than with PPO. This advantage stems from GRPO’s rollout mechanism, which enables i-MENTOR to generate varied exploration rewards for multiple responses to the same question. Such a design promotes broader exploration of potential solution paths compared to single-response PPO updates.
3.3 Analysis on Multiple Generations (Q1)
In this section, we validate i-MENTOR’s efficacy on the AIME 2024 benchmark between GRPO and i-MENTOR-GRPO. For each question, the LLM generates multiple answers to enable more rigorous performance evaluation (Yu et al., 2025). Results in Table 2 demonstrate two key findings: First, LLM accuracy improves consistently with increasing response counts, confirming that aggregating multiple outputs enhances model performance. Second, i-MENTOR achieves approximately 20% improvement across all metrics compared to GRPO baselines. This validates i-MENTOR’s effectiveness in solving complex problems while efficiently leveraging multiple responses.
@@ 3.4 Generalization Analysis (Q2)
In this section, we conduct experimental analysis to verify the generalization ability of i-MENTOR from both the model and dataset perspectives.
@@ 3.4.1 Base Model Generalization Analysis
Here, we verify the effectiveness of i-MENTOR on a new commonly used base model, deepseek-7b, with GRPO on the GSM8K dataset to validate i-MENTOR’s generalization on different base models. From the result in Table 3, we could conclude that i-MENTOR can be applied to different base models, including Qwen and DeepSeek, and provide performance gains for their reasoning process.
Table 3: Experimental result with deepseek-7b base model on GSM8K.
Dataset GRPO i-MENTOR-GRPO
GSM8K 0.6535 0.6641
@@ 3.4.2 Dataset Generalization Analysis
This subsection examines i-MENTOR’s impact on generalization capability by evaluating performance on out-of-distribution data. Specifically, we train the LLM exclusively on Countdown-4 and measure performance on the unseen Countdown-34 dataset. As evidenced by Table 4, i-MENTOR-GRPO surpasses GRPO on both Countdown-4 and the held-out Countdown-34 dataset after training. This demonstrates that i-MENTOR not only enhances optimization efficacy but also preserves - and even strengthens - the model’s generalization capacity when exposed to novel data distributions.
@@
Dataset GRPO iMENTOR-GRPO Improve
Coundown-4 0.3831 0.4726 23.36%
Countdown-34 0.3369 0.3877 15.08%
@@
Dataset PPO i-MENTOR-PPO GRPO i-MENTOR-GRPO
@@ 4 Training Time Analysis
Here we summarize the training time for experiments in Table 1 to examine the additional cost in terms of time that i-MENTOR incurs. The results in Table 5 demonstrate that i-MENTOR introduces only minimal training time overhead versus baselines without it. This efficiency stems from the predictor and target networks requiring small parameter counts–they comprise several compact linear layers. These networks compute rewards via output differences to determine trajectory visit frequency during training, without needing complex semantic understanding. Critically, since the exploration module operates solely during training, i-MENTOR has zero inference overhead. This confirms i-MENTOR’s superiority in delivering performance gains with negligible training time impact.
@@ 5.1 Reinforcement Learning for LLMs
Reinforcement learning has evolved from foundational approaches like PPO (Schulman et al., 2017) to advanced frameworks such as GRPO (Shao et al., 2024). PPO achieves stability in policy updates through clipped objectives and trust region constraints, effectively aligning models for dialogue systems and code generation (Wang et al., 2024b; Shojaee et al., 2023). Though joint policy-value optimization introduces computational overhead (Shao et al., 2024). GRPO (Shao et al., 2024) mitigates these limitations through trajectory-level sampling and group-wise advantage normalization. By generating multiple responses per input and standardizing rewards across trajectory groups, GRPO reduces policy update bias while maintaining efficiency. Recent implementations such as DAPO (Yu et al., 2025) and Dr. GRPO (Liu et al., 2025) further showcase GRPO’s scalability through detailed refinement of reasoning objectives, though exploration efficiency remains constrained by static and sparse reward structures (Dou et al., 2025; Gao et al., 2024). Differently, i-MENTOR enhances LLM reasoning through providing dense, stable, and efficient exploration rewards in RL optimization process. These rewards enable intrinsic motivation-guided exploration over diverse response trajectories during training, systematically improving reasoning capabilities.
@@ 5.2 Improving Reasoning Ability of LLMs
Recent advances in LLM reasoning focus on three key paradigms: Pre-training augmentation improves foundational reasoning by exposing models to curated datasets and synthetic traces. Models like Qwen (Yang et al., 2024b) and Llama (Grattafiori et al., 2024) show significant gains through domain-specific data scaling, though this requires heavy computational resources and careful data curation. Prompt engineering (Sahoo et al., 2024; Chen et al., 2023) enhances reasoning via structured inputs. Techniques like Chain-of-Thought (Yu et al., 2023; Chu et al., 2023) decompose problems into steps, while self-consistency (Wang et al., 2022) improves outputs through majority voting. Extensions such as multi-agent debates (Liang et al., 2023) and iterative refinement (Madaan et al., 2023) further extend capabilities but are limited by manual design and high computational costs. Algorithmic enhancement combines inference-time search with RL-based fine-tuning. Methods like Monte Carlo Tree Search (Browne et al., 2012) and RL approaches (e.g., PPO and GRPO) optimize reasoning policies but often overexploit fixed high-reward paths, limiting exploration of novel solutions. Unlike these methods, i-MENTOR actively incentivizes exploration of under-optimized paths through exploration rewards, enabling deeper action-space traversal.
@@ 6 Conclusion
We introduce i-MENTOR, a reinforcement learning framework that systematically addresses the challenge of sparse rewards in LLM reasoning through three methodological advances. First, our trajectory-aware exploration reward employs lightweight networks to provide dense intrinsic rewards while eliminating computational overhead and episodic length bias, enabling efficient discovery of high-quality reasoning paths. Second, we propose error-conditioned reward allocation to ensure efficient exploration on challenging samples while intrinsically stabilizing training in large action spaces, ensuring sustained exploration. Third, our advantage-preserving integration decouples exploration incentives from outcome-based policy gradients, resolving fundamental conflicts between intrinsic and extrinsic rewards. Extensive experiments demonstrate that i-MENTOR significantly enhances reasoning performance when integrated with both PPO and GRPO, establishing its versatility across distinct RL paradigms.

### v6

@@ 2.2.2 Error-conditioned Reward Allocation
Despite the benefits of trajectory-aware exploration rewards, the large action space in LLM reasoning still introduces training challenges. Naive exploration strategies become infeasible due to unstable and sparse rewards and suboptimal exploration efficiency. Moreover, uniform exploration rewards across correct and incorrect samples reduce exploration efficiency and induce instability in correct trajectories. To enable more effective exploration, we prioritize samples with residual errors to facilitate diverse discoveries while efficiently improving overall performance. Thus, IMAGINE employs an error-conditioned reward allocation mechanism across sampling sequences. Specifically, a selection function $I_{a\neq o}$ is designed to allocate exploration rewards exclusively to incorrect samples:
@@ 2.2.3 Advantage-preserving Integration
Since the intensity of the exploration behavior varies at different stages of training, directly combining exploration rewards with outcome-based rewards creates conflicting learning signals across training stages. For PPO, this injects noise into value function estimation. For GRPO, exploration rewards may invert originally positive advantage signs after group normalization, which is a critical issue since exploration rewards should never penalize samples. To resolve this conflict and ensure seamless integration with established RL algorithms like PPO and GRPO, IMAGINE applies exploration rewards after advantage computation and possible value estimation through:
@@
where $\hat{A}_{\text{old}}$ denotes the original advantage from outcome-based rewards. For PPO, we apply the exploration reward to the last advantage token. For GRPO, we apply it to each advantage in a rollout. This design preserves two key properties: (1) The outcome-based advantage maintains its original mean and variance statistics, ensuring stable policy updates; (2) Exploration rewards $R^{\star}(q,o)$ provide trajectory-level incentive-based guidance without distorting value estimation. The decoupled formulation prevents exploration rewards from conflicting with outcome-based advantage normalization in GRPO or perturbing value function update in PPO, enabling harmonious integration of both reward components.
By applying trajectory-aware exploration rewards with error-conditioned reward allocation and advantage-preserving integration, IMAGINE efficiently enhances LLM reasoning by guiding the model toward diverse possible responses during training, while significantly improving overall reasoning ability. Algorithmically, IMAGINE employs lightweight networks that operate independently of the LLM, enabling efficient updates and modular reward computation. This design allows IMAGINE to integrate seamlessly with standard RLVR algorithms such as PPO and GRPO, consistently yielding performance improvements. The complete algorithm is detailed in Appendix B.
@@ 3 Experiments
Here we evaluate IMAGINE on three datasets to investigate the following research questions:
  • Q1: How does IMAGINE enhance LLM reasoning performance?
  • Q2: How does IMAGINE generalize across models and data distributions?
In the following subsections, we begin by outlining the experimental setup, followed by a systematic results analysis addressing our core questions. In addition, due to limited space, Appendix A also provides an expanded case study for Figure 2, while implementation specifics–including pseudo code (Appendix B), data statistics (Appendix C), and configuration details (Appendix D)–are deferred for completeness. Beyond core experiments, Appendix E validates IMAGINE’s superiority over basic exploration techniques. We further include an ablation study (Appendix F) quantifying submodule contributions, a sensitivity analysis (Appendix G) examining hyperparameter impacts on performance, and a group size analysis in Appendix H for robustness validation.
@@ 3.1.1 Dataset
Following previous studies (Yu et al., 2025; Pan et al., 2025), we conduct experiments on three public datasets, i.e., GSM8K(Cobbe et al., 2021), Countdown-34 and its harder version Countdown-4(Gandhi et al., 2024) to validate IMAGINE’s effectiveness against vanilla PPO and GRPO algorithms. For computational efficiency, we use a subset of the complete dataset of Countdown-34 and Countdown-4 for training. The detailed dataset statistic is shown in Appendix C. Additionally, we also provide detailed experiments on the AIME 2024 dataset between GRPO and IMAGINE (GRPO) to further validate IMAGINE’s effectiveness with multiple outputs in Table 2.
Table 1: Average accuracy. IMAGINE (PPO)/(GRPO) denotes the implementation of IMAGINE on PPO/GRPO. “Improve” indicates relative improvement.
Dataset PPO IMAGINE (PPO) Improve GRPO IMAGINE (GRPO) Improve
[…]
@@
3.1.2 Evaluation Protocol
Network structure of IMAGINE: The predictor and target networks in IMAGINE share the same architecture, which is dataset-agnostic and LLM-independent. For computational efficiency, we simply apply a lightweight predictor and target network structure for IMAGINE (embedding size=16=16, three FFN layers with neurons $[16,8,1]$), ensuring negligible additional training time. To make a fairer comparison, we apply a set of fixed training hyperparameters to all datasets.
3.2 Main Result (Q1)
The main results are presented in Table 1. Figure 5 highlights key training details via some training trajectories, showcasing the performance improvements achieved during training. From Table 1 and Figure 5 we could conclude that:
  • GRPO outperforms PPO on all datasets, illustrating superior training effects brought by group-wise sampling and reward normalization. Both IMAGINE (PPO) and IMAGINE (GRPO) outperform standard RL approaches, demonstrating that IMAGINE effectively guides and enhances the updates of LLM reasoning via exploration behavior during the RL training process. By encouraging the exploration of new responses rather than merely fitting to outcome-based reward through sampling behaviors, IMAGINE enables LLMs to explore more diverse potential inference paths during training and avoids LLMs getting trapped in local optimal solutions.
  • A key challenge in improving LLM reasoning lies in handling difficult samples, which typically yield near-zero rewards that hinder parameter updates. Our experiments reveal an interesting pattern: IMAGINE’s improvements are most significant on Countdown-4, followed by Countdown-34 and GSM8K. This progression aligns with the relative difficulty levels of these datasets, where Countdown-4 > Countdown-34 > GSM8K. The results suggest that IMAGINE’s dense exploration rewards specifically help models overcome learning barriers in challenging samples. This phenomenon represents a further contribution of our study: it originates from RLVR’s core distinction from traditional RL–leveraging a pre-trained LLM instead of a randomly initialized model. Consequently, exploration yields significantly greater benefits when addressing complex problems.
  • IMAGINE achieves greater performance gains with GRPO than with PPO. This advantage stems from GRPO’s rollout mechanism, which enables IMAGINE to generate varied exploration rewards for multiple responses to the same question. Such a design promotes broader exploration of potential solution paths compared to single-response PPO updates.
@@ 3.2 Main Result (Q1)
3.3 Analysis on Multiple Generations (Q1)
In this section, we validate IMAGINE’s efficacy on the AIME 2024 benchmark between GRPO and IMAGINE (GRPO) following previous studies (Yu et al., 2025; Liu et al., 2025). For each question, the LLM generates multiple answers to enable more rigorous performance evaluation (Yu et al., 2025). Results in Table 2 demonstrate two key findings: First, LLM accuracy improves consistently with increasing response counts, confirming that aggregating multiple outputs enhances model performance. Second, IMAGINE achieves approximately 20% improvement across all metrics compared to GRPO baselines. This validates IMAGINE’s effectiveness in solving complex problems while efficiently leveraging multiple responses.
@@ 3.4 Generalization Analysis (Q2)
In this section, we conduct experimental analysis to verify the generalization ability of IMAGINE from both the model and dataset perspectives.
@@ 3.4.1 Base Model Generalization Analysis
Here, we verify the effectiveness of IMAGINE on another commonly used base model, deepseek-7b, with GRPO on the GSM8K dataset to validate IMAGINE’s generalization on different base models. From the result in Table 3, we could conclude that IMAGINE can be applied to different base models, including Qwen and DeepSeek, and provide stable performance gains for their reasoning process.
Table 3: Experiments with deepseek-7b base model on GSM8K.
Dataset GRPO IMAGINE (GRPO)
[…]
@@ 3.4.2 Dataset Generalization Analysis
This subsection examines IMAGINE’s impact on generalization capability by evaluating performance on out-of-distribution data. Specifically, we train the LLM exclusively on Countdown-4 and measure performance on the unseen Countdown-34 dataset. As evidenced by Table 4, IMAGINE (GRPO) surpasses GRPO on both Countdown-4 and the held-out Countdown-34 dataset after training. This demonstrates that IMAGINE not only enhances optimization efficacy but also preserves - and even strengthens - the model’s generalization capacity when exposed to novel data distributions.
@@
Dataset GRPO IMAGINE (GRPO) Improve
[…]
@@
Dataset PPO IMAGINE (PPO) GRPO IMAGINE (GRPO)
@@ 4 Training Time Analysis
Here we summarize the training time for experiments in Table 1 to examine the additional cost in terms of time that IMAGINE incurs. The results in Table 5 demonstrate that IMAGINE introduces only minimal training time overhead versus baselines without it. This efficiency stems from the predictor and target networks requiring small parameter counts–they comprise several compact linear layers. These networks compute rewards via output differences to determine trajectory visit frequency during training, without needing complex semantic understanding. Critically, since the exploration module operates solely during training, IMAGINE has zero inference overhead. This confirms IMAGINE’s superiority in delivering performance gains with negligible training time impact.
5 Exploration Analysis
In this section, we quantified the exploration effect of GRPO and IMAGINE (GRPO) during training to analyze the promoting role of IMAGINE on exploration behavior. Specifically, as Section 2.2.1 and Equation (5) have stated, the convergence rate of the loss $\mathcal{L}(q,o)$ provided by IMAGINE represents the frequency of occurrence of similar trajectories during the training process. The higher the frequency of similar trajectories, the faster $\mathcal{L}(q,o)$ decreases on this part of the trajectory, and at the same time, the smaller the average value of $\mathcal{L}(q,o)$ within such a batch. Therefore, the decrease speed of this value can be used as a criterion to judge the diversity of trajectories in the training stage.
Figure 6: Exploration analysis on Countdown-4.
As in Figure 6, the decrease speed of $\mathcal{L}(q,o)$ on GRPO is significantly faster than that of IMAGINE (GRPO). This shows that the trajectories generated during training with IMAGINE (GRPO) are more diverse than those with GRPO.
6 Scalability Analysis
To further test the scalability of IMAGINE on different model sizes, we experiment the performance of IMAGINE (GRPO) on Countdown-34 with different base model sizes: Qwen2.5- 0.5B/1.5B/3B. The results are shown in Table 6. From the table, we can conclude that IMAGINE can be applied to models of different sizes, bringing stable performance improvements.
Table 6: Scalability analysis on Countdown-34.
Model GRPO IMAGINE (GRPO)
Qwen2.5-0.5B 0.0033 0.0107
Qwen2.5-1.5B 0.4532 0.5068
Qwen2.5-3B 0.6711 0.7132
7 Compatibility Analysis
Owing to IMAGINE ’s lightweight structure and strong compatibility, it can be applied directly to PPO and GRPO, and it can also be combined with recent advanced RL variants such as DAPO (Yu et al., 2025), potentially facilitating exploration and improving performance. In this section, we further examine IMAGINE ’s compatibility with GRPO-based techniques. Accordingly, we evaluate the combination of DAPO and IMAGINE on Countdown-4. As shown in Table 7, the combined method consistently outperforms either component alone, suggesting that IMAGINE is complementary to DAPO and supporting its applicability across different training variants.
Table 7: Compatibility analysis.
Method Accuracy
DAPO 0.4691
IMAGINE (DAPO) 0.4924
@@ 8.1 Reinforcement Learning for LLMs
Reinforcement learning has evolved from foundational approaches like PPO (Schulman et al., 2017) to advanced frameworks such as GRPO (Shao et al., 2024). PPO achieves stability in policy updates through clipped objectives and trust region constraints, effectively aligning models for dialogue systems and code generation (Wang et al., 2024b; Shojaee et al., 2023). Though joint policy-value optimization introduces computational overhead (Shao et al., 2024). GRPO (Shao et al., 2024) mitigates these limitations through trajectory-level sampling and group-wise advantage normalization. By generating multiple responses per input and standardizing rewards across trajectory groups, GRPO reduces policy update bias while maintaining efficiency. Recent implementations such as DAPO (Yu et al., 2025) and Dr. GRPO (Liu et al., 2025) further showcase GRPO’s scalability through detailed refinement of reasoning objectives, though exploration efficiency remains constrained by static and sparse reward structures (Dou et al., 2025; Gao et al., 2024). Differently, IMAGINE enhances LLM reasoning through providing dense, stable, and efficient exploration rewards in RL optimization process. These rewards enable intrinsic motivation-guided exploration over diverse response trajectories during training, systematically improving reasoning capabilities.
@@ 8.2 Improving Reasoning Ability of LLMs
Recent advances in LLM reasoning focus on three key paradigms: Pre-training augmentation improves foundational reasoning by exposing models to curated datasets and synthetic traces. Models like Qwen (Yang et al., 2024b) and Llama (Grattafiori et al., 2024) show significant gains through domain-specific data scaling, though this requires heavy computational resources and careful data curation. Prompt engineering (Sahoo et al., 2024; Chen et al., 2023) enhances reasoning via structured inputs. Techniques like Chain-of-Thought (Yu et al., 2023; Chu et al., 2023) decompose problems into steps, while self-consistency (Wang et al., 2022) improves outputs through majority voting. Extensions such as multi-agent debates (Liang et al., 2023) and iterative refinement (Madaan et al., 2023) further extend capabilities but are limited by manual design and high computational costs. Algorithmic enhancement combines inference-time search with RL-based fine-tuning. Methods like Monte Carlo Tree Search (Browne et al., 2012) and RL approaches (e.g., PPO and GRPO) optimize reasoning policies but often overexploit fixed high-reward paths, limiting exploration of novel solutions. Unlike these methods, IMAGINE actively incentivizes exploration of under-optimized paths through exploration rewards, enabling deeper action-space traversal.
@@ 9 Conclusion
We introduce IMAGINE, a reinforcement learning framework that systematically addresses the challenges of sparse rewards and limited exploration in RL-based LLM finetuning through three methodological advances. First, our trajectory-aware exploration reward employs lightweight networks to provide dense exploration rewards, enabling efficient discovery of high-quality reasoning paths. Second, we propose error-conditioned reward allocation to ensure efficient exploration on challenging samples while intrinsically stabilizing training in large action spaces, ensuring sustained exploration. Third, our advantage-preserving integration decouples exploration incentives from outcome-based policy gradients, resolving fundamental conflicts between intrinsic and extrinsic rewards. Extensive experiments demonstrate that IMAGINE significantly enhances reasoning performance when integrated with both PPO and GRPO, establishing its versatility across distinct RL paradigms.